Tally Corp. sells software during the recruiting seasons. During the current​ year,

10,000

software packages were sold resulting in

$470,000

of sales​ revenue,

$130,000

of variable​ costs, and

$48,000

of fixed costs.If sales increase by

$80,000​,

operating income will increase by​ ________. (Round interim calculations to two decimal places and the final answer to the nearest whole​ dollar.)

A.$48,000

B.$30,588

C.$57,872

D.$32,000

While−You−Train

is a

not−for−profit

organization that aids the unemployed by supplementing their incomes by​ $7,000 annually, while they seek new employment skills. The organization has fixed costs of​ $230,000 and the budgeted appropriation for the year totals​ $720,000. How many individuals can receive financial assistance this​ year?

A.70 people

B.103 people

C.136 people

D.33 people

A manufacturing plant produces two product​ lines: golf equipment and soccer equipment. An example of indirect cost for the soccer equipment line is the​ ________.

A.material used to manufacture the soccer studs

B.labor to shape the leather used to make the soccer ball

C.material used to make the soccer balls

D.property taxes paid on the land and building​ (plant)
Answer #1

1.

Contribution margin = $470,000 - 130,000 = $340,000

Contribution margin ratio = $340,000 / 470,000 = 72.34%

Operating income will increase by = $80,000 * 72.34% = $57,872

Hence option C is correct.

2.

Individual to receive financial assistance = $720,000 - $230,000 / $7,000

Individual to receive financial assistance = $490,000 / $7,000 = 70 people

Hence option A is correct

3.

Property taxes paid on the land and building (plant)

Hence option D is correct.